OpenStudy (anonymous):

Does anyone know what is a simplified form of the expression log21(5) +log21(4) -log21(2)? All I know is that the answer is log21 (10), but I don't know how to get that answer.

ganeshie8 (ganeshie8):

use below :- \(\log a + \log b = \log (ab)\) \(\log a - \log b = \log (a/b)\)

OpenStudy (anonymous):

wait, is log a log 21(5) and log b log 21(4)?

ganeshie8 (ganeshie8):

\(\log_{21}(5) +\log_{21}(4) -\log_{21}(2)\)

ganeshie8 (ganeshie8):

\(\log_{21}(5\times 4) -\log_{21}(2)\)

ganeshie8 (ganeshie8):

\(\log_{21}(5\times 4/2)\)

ganeshie8 (ganeshie8):

\(\log_{21}(20/2)\)

ganeshie8 (ganeshie8):

\(\log_{21}(10)\)

OpenStudy (anonymous):

oh ok, thank-you

OpenStudy (anonymous):

so, do we always use that formula to solve a problem like that?

ganeshie8 (ganeshie8):

yes, u can always use that formula as long u have the same base...

ganeshie8 (ganeshie8):

base is 21 for our current problem

OpenStudy (anonymous):

oh, ok thank-you

OpenStudy (anonymous):

again

ganeshie8 (ganeshie8):

np :)